# Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice

### v3

@@ Abstract
Social emotions such as shame and pride reflect social sanctions or approvals in society. In this paper, we examine how expressions of shame and pride vary across cultures and harness them to extract unspoken normative expectations across cultures. We introduce the first cross-cultural shame/pride emotions movie dialogue dataset, obtained from  5.4K Bollywood and Hollywood movies, along with over $10K$ implicit social norms. Our study reveals variations in expressions of social emotions and social norms that align with known cultural tendencies observed in the United States and India -- e.g., Hollywood movies express shame predominantly toward self whereas Bollywood movies express shame predominantly toward others. Similarly, Bollywood shames non-conformity in gender roles, and takes pride in collective identity, while Hollywood shames lack of accountability, and takes pride in ethical behavior. More importantly, women face more prejudice across cultures and are sanctioned for similar social norms.
@@ 1 Introduction
Normative expectations [^1] Bicchieri (2005) uses the term normative expectations, Lahti and Weinstein (2005) use moral norms, and many social psychologists Cialdini et al. (1990) use injunctive norms. We use the terms social norms and normative expectations interchangeably., a category of social norms that refer to others’ beliefs about how we ought to behave Bicchieri (2005). It reveals attitudes and values that influence social actions across cultures. For instance, hesitancy toward seeking mental help could exist because others expect an individual not to seek help or because the individual believes that others do not seek help. The former reveals the attitude of society, whereas the latter indicates an individual’ lack of awareness. Capturing diversity in cultural beliefs is paramount to building socially aligned AI models for diverse users Talat et al. (2021); Atari et al. (2023), for inducing collective behavior change (e.g., overcoming stigma against mental health disorders) and several other social applications.
@@
Prior works crowd-sourced social norms from Q&A style social media posts (such as Reddit) Jiang et al. (2021); Nahian et al. (2020) or prompted large language models (LLMs) Fung et al. (2022) without distinguishing between normative and empirical expectations. However, social media platforms [^2] Twitter Statistic: https://blog.hootsuite.com/twitter-statistics/, [^3] Reddit Statistic: https://www.enterpriseappstoday.com/stats/reddit-statistics.html are skewed toward Western demographics (i.e., young white male) and the reliance on a small cohort of annotators typically located in Western countries further enforces Anglocentric beliefs when collecting socially acceptable behaviors Ziems et al. (2022). LLMs with sophisticated capabilities such as GPT-4 are increasingly used to extract social norms; however, Havaldar et al. (2023a) illustrated their lack of cross-cultural understanding.
@@
The overarching question we ask in this paper is, how and what social behaviors are sanctioned or encouraged across cultures. Shame and pride are social emotions experienced across cultures on violating or adhering to normative expectations in society Tracy and Robins (2004). Notably, the function of shame (i.e., social regulation) is consistent across cultures Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022). Examining situations evoking shame and pride can thus reveal unspoken social norms specific to a culture (See Table 1).
In this paper, we (a) examine cross-cultural variations in linguistic expressions of shame and pride, and (b) propose a cross-cultural social emotion-based approach for norm discovery. The appraisals and expressions of shame and pride vary across cultures and is a research question of great significance to understand society and self-regulation Wong and Tsai (2007); Lewis et al. (2010); Wegge et al. (2022). We chose two top-grossing movie industries, i.e., Hollywood and Bollywood, to analyze how social emotions and norms differ across cultures. Hollywood primarily depicts social situations from the USA, an individualist society that values competency and autonomy Triandis (1989, 1988) whereas Bollywood depicts India, a collectivist society where one’s sense of self is interwoven with community beliefs. The cultural dichotomy between India and the USA (i.e., collectivism vs. individualism) thus presents a rich ground for understanding variations in their beliefs and values.
We release a large cross-cultural dataset of social emotions-related dialogues with $10K$ social norms. Using a blend of psychology-informed language analysis and LLM-powered approach to mine implicit information, we demonstrate cross-cultural variations in how and why these emotions are expressed. In parallel, we examine how social sanctions are attributed to men and women across cultures influencing their social behaviors; women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021). We hypothesize that women are more subjected to social sanctions than men across cultures.
2 Movies Subtitles Corpora
Movies provide life-like depictions of social situations to study how and why these emotions are expressed. They are influenced by and/or, known to influence real lives Adkins and Castle (2014); Kubrak (2020). In contrast, social media Q&A posts are (a) skewed to social situations specific to tech-savvy young demographic groups, and (b) are second-hand reports that do not reveal fully the social dynamics of how these emotions are expressed in natural conversations, for instance, between a boss and an employee or among family members at a Christmas table.
For our study, we collected English subtitles for 5,435 Hollywood and Bollywood movies that were released post-1990 by auto-crawling websites that host or link movie subtitles (See Table A2 for data distribution). The year of release for movies was verified by either parsing subtitle file names having a release year or checking Wikipedia entries. The year mapping was performed to ensure a similar period for collected movies.
2.1 Extracting Shame and Pride related Expressions
The expressions of shame and pride could be implicit (as indicated by lowered shoulders, avoiding eye contact, etc.), and explicit (See Table 1). Detecting implicit social emotions in textual discourse is an innately complex task requiring socio-cultural knowledge that may not be evident to outsiders or third-party annotators. Below is one such example:
‘‘Actor 1: Liza, Alice! Come away from that window this minute. I think it’s disgusting, standing there where they can see you.
Actors 2,3 (Liza, Alice): I notice that you peek out often enough when you think no one’s looking.
Actor 1: I never!
Actors 2,3 (Liza, Alice): I saw you! So did I.
Here, it is difficult to determine if shame is evoked and/or experienced by actors in this discourse. A third-party annotation will reveal their beliefs and attitudes, which could introduce bias in our analysis. Adopting a high precision-low recall strategy, we focus on explicit mentions of shame and pride to learn about occurrences of shame and pride as determined by the characters. We extracted dialogues having the word shame or pride or their variations (See Appendix A1), along with the previous and the next two lines for situational context (See Table 1). In the case of short dialogues, which could be the case due to monosyllabic responses in spoken conversations, an extra previous and next line was appended to context.
Manual Screening One author manually checked the entire dataset of shame and pride-related dialogues to filter out dialogues unrelated to social sanctions/approvals e.g., phrases such as "what a shame", "it’s a shame", "proudly presents". Below is an example of such dialogue.
‘‘…His music is sad all of the time. It’s really a shame. He’ll never be allowed to go up to heaven.   But you will, won’t you, Johan?   Oh yes. I’ll get there."
Similarly, the phrase "proudly presents" is used while introducing an event or an artist in Hollywood movies.
Final Dataset: We created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood. Cross-cultural differences also exist in speaking styles across cultures and we thus formed a control set of dialogues unrelated to shame and pride for both movie industries to remove the variations in language markers owing to culture-specific speaking styles (See Table A2 for dialogue distribution).
@@ 2.1 Extracting Shame and Pride related Expressions
Table 2: Pearson r for LIWC categories significantly correlated ($p<0.05$) with shame for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values). Note the contrast that is, Self-centered shame vs Other-centered shame; Past focus vs Present focus in Hollywood vs Bollywood. See Table A7 and A9 for the complete set of correlations.
@@ 3 Approach
Fig. 1 illustrates our study design comprising a blend of two key approaches (a) a psychosocial vocabulary approach to measure how expressions of shame and pride differ and (b) prompting Large Language Model approach to extract reasons behind them and how they differ between India and the U.S.A.
@@ 3.2 Pretrained LLM: GPT-4 Chat
LLMs such as GPT-4 exhibit superlative English language understanding and are more reliable than human annotators on platforms such as MTurk Pan et al. (2023); Törnberg (2023). We used GPT-4 chat in two-shot setting ( See Tables A5 and A6 for prompts) to seek answers to the following questions:
  • who is $<$being shamed/feeling proud$>$ in the given movie discourse, and what is their gender?
@@
Since there are at least two characters in a discourse, the first question orients the LLM to focus on the person experiencing the social sanction or approval and then identify their gender. The output for the second question serves as the implicit social norm in the culture. Asking for "reason" leading to the expression of social emotion encourages LLMs to retrieve norm from provided context, mitigating potential Anglocentric tendencies.
Thematic Analysis of Social Norms
To capture overarching themes in reasons behind expressing shame and pride in Indian and American societies, we performed agglomerative clustering after embedding unique shame and pride-related social norms using SBERT embeddings Reimers and Gurevych (2019). Unlike movie dialogues, social norms extracted from GPT-4 chat are short phrases (See Table 4) and are devoid of culture-specific language style markers (i.e., Indian English vs American English) thus, no control for culture is needed.
@@ 4 Results
4.1 How do expressions of shame/pride vary across cultures?
Shame is associated with negative emotions, power, and morality in both movie industries (See Table 2), however, significant cultural nuances exist in its manifestation. In Hollywood, shame is more I-focused with remorse (e.g., sadness, anxiety) and past-focused language, whereas in Bollywood, it is more You-focused with frequent social references and present-focused language. In Bollywood, psycho-social categories associated with shame reveal a larger social context (note the contrast in family i.e., nuclear family in Hollywood vs extended family in Bollywood) and social situations (i.e., anger, swear, conflict). Categories female and sexual are exclusively correlated with shame in Bollywood potentially indicating an honor system in collectivist societies Caffaro et al. (2014) whereas it is illness that is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
A similar dichotomy is observed for Pride related interactions (See Table 3). Pride in Hollywood interactions is more centered on family (note, family in Bollywood is male-centric and comprises extended family (such as uncle)) and self-centered (e.g., I am proud of __) whereas it is achievement/power focused and "We"-centered in Bollywood movies (with othering (us vs they)- See Appendix). Male references are more strongly associated with pride than females in Hollywood whereas the female category has an insignificant ($p>0.05$) correlation with pride in Bollywood. Ethnicity-related pride is expressed in both movie industries but political category is exclusively correlated to Bollywood. Pride is more positive (see POSEMO in Table 3) in Hollywood whereas in Bollywood, it reflects honor (e.g., protecting family’s pride, bringing pride to family).
@@ 4.1 How do expressions of shame/pride vary across cultures?
4.2 How do reasons behind shame/pride vary across cultures?
Cross-Cultural Norms Dataset
For the Bollywood set, GPT-4 chat predicted reasons (also implicit social norms) for 5321 (98.4%) shame-related dialogues out of 5409, and 2237 (74.6%) pride-related dialogues out of 2999. For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805. Upon manual analysis, we found that pride is also used to express affection, specifically toward close family members such as sons - therefore, a lower number of norms is associated with pride compared to shame. GPT-4 chat predicted a total of 10,445 social norms (See Table A3 for norm distribution). Prompting "reasons" behind shame and pride allowed us to capture high specificity in cultural norms (See Tables 4 and 6 for examples).
4.2.1 Manual Evaluation
Two volunteers manually verified the predicted gender for "the person experiencing shame/pride" and the predicted "reason" in a randomly sampled set of $100$ dialogues each from Hollywood and Bollywood. The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society. Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social norms of the U.S.A.

### v4

@@ Abstract
Shame and pride are social emotions expressed across cultures to motivate and regulate people’s thoughts, feelings, and behaviors. In this paper, we introduce the first cross-cultural dataset of over $10k$ shame/pride-related expressions, with underlying social expectations from  5.4K Bollywood and Hollywood movies. We examine how and why shame and pride are expressed across cultures using a blend of psychology-informed language analysis combined with large language models. We find significant cross-cultural differences in shame and pride expression aligning with known cultural tendencies of the USA and India -- e.g., in Hollywood, shame-expressions predominantly discuss self whereas Bollywood discusses shame toward others. Pride in Hollywood is individualistic with more self-referential singular pronouns such as I and my whereas in Bollywood, pride is collective with higher use of self-referential plural pronouns such as we and our. Lastly, women are more sanctioned across cultures and for violating similar social expectations e.g. promiscuity.
@@ 1 Introduction
Shame and pride are social emotions expressed across cultures to motivate and regulate people’s thoughts, feelings, behaviors, achievements, and social adaptations Fischer and Tangney (1995); Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022). The expressions and appraisals of shame and pride vary across cultures and is a research question of great significance to understand society and self-regulation Wong and Tsai (2007); Lewis et al. (2010); Wegge et al. (2022). The overarching question we ask in this paper is, how and why shame and pride are expressed across cultures. Knowing cross-cultural differences in shame/pride-related expressions will help us measure psychological constructs such as shame proneness Tangney et al. (1992); Cohen et al. (2011) and self-esteem Pyszczynski et al. (2004) reliably from language -- enabling effective human-computer interactions in domains such as AI-driven psychotherapy Resendiz and Klinger (2023); Demszky et al. (2020). Examining behaviors evoking shame and pride will reveal culture-specific societal expectations and attitudes (See Table 1) -- paramount to building culturally competent AI models for diverse users Talat et al. (2021); Atari et al. (2023).
@@
For our study, we select two top-grossing movie industries, i.e., Hollywood and Bollywood  [^1] https://nairametrics.com/2024/08/14/top10-fastest-growing-film-industries-in-the-world/. Hollywood primarily depicts social situations from the USA, an individualist society that values competency and autonomy Triandis (1989, 1988). Bollywood depicts India, a collectivist society where one’s sense of self is interwoven with community beliefs. The cultural dichotomy between India and the USA (i.e., collectivism vs. individualism) thus presents a rich ground for understanding variations in their beliefs and values. Our approach blends psychology-informed language analysis with state-of-the-art Large Language Model (LLM) to delineate interpretable psychosocial constructs such as self-focus, morality, etc. in shame/pride-related expressions and extract implicit social expectations behind them.
This paper makes the following research contributions: (a) We build a cross-cultural dataset of over 10k shame/pride-related dialogues with underlying social expectations  [^2] Link will be provided upon acceptance, (b) We demonstrate significant cross-cultural variations in how and why shame and pride are expressed in movies, and (c) We illustrate that women are more subjected to social sanctions than men across cultures and for violating similar social expectations.
To the best of our knowledge, this is the first empirical study analyzing the psycho-social constructs underlying shame/pride-related expressions across cultures and the social expectations behind them.
2 Movies Subtitles Corpora
Movies provide culture-specific life-like depictions of social situations Adkins and Castle (2014); Kubrak (2020). The natural conversation style between characters in movies can reveal the social power dynamics (e.g., boss-employee, father-daughter) and gender roles. Another potential alternative is social media posts however, they are (a) skewed to social situations specific to young demographic groups from Western countries, and (b) are second-hand reports that do not reveal fully the social dynamics of how shame/pride are expressed in natural conversations, for instance, between a boss and an employee or among family members at a Christmas table.
We collected English subtitles for 5,435 Hollywood and Bollywood movies that were released post-1990 by auto-crawling websites that host or link movie subtitles (See Table 2 for data distribution). The subtitles are professionally done translations to preserve the emotions and intended message. Bollywood movie subtitles were collected from www.Bollynook.com and Hollywood movie subtitles were from the publicly available Kaggle movie subtitles dataset [^3] https://www.kaggle.com/datasets/adiamaan/movie-subtitle-dataset. The year of release for movies was verified by either parsing subtitle file names having a release year or checking Wikipedia entries. The year mapping was performed to ensure a similar period for collected movies.
Hollywood Bollywood
#Movies 2697 2738
#Tokens 20.78M 22.62M
#Shame 1221 5409
#Pride 2805 2999
#Control 4385 8303
Table 2: Data Distribution. #shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless - See Table A1). Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud). Control is the group of dialogues without words shame, pride, and their derivatives.
2.1 Extracting Shame and Pride related Expressions
The expressions of shame and pride could be explicit (See Table 1) or implicit (e.g. characters’ lowered shoulders, avoiding eye contact, etc.). We extracted explicit mentions of shame and pride to learn about occurrences of shame and pride as determined by the characters. We adopted a keyword search-based approach (See Appendix A1) to identify dialogues along with the previous and the next two lines for situational context (See Table 1). In the case of short dialogues, which could be the case due to monosyllabic responses in spoken conversations, an extra previous and next line was appended to context. One author manually checked the entire dataset of shame and pride-related dialogues to filter out dialogues with phrases such as "what a shame", "it’s a shame", "proudly presents".
Detecting implicit shame/pride-related expressions in textual discourse is an innately complex task. Computational systems for detecting social emotions such as shame and guilt underperform significantly compared to basic emotions such as joy Demszky et al. (2020); Resendiz and Klinger (2023) due to internal self-evaluation associated with social emotions. We tested the GPT-4 language model Achiam et al. (2023) to detect implicit expressions of shame in given movie dialogues without explicit mentions of shame and pride. We recruited two human annotators to determine the agreement with GPT-4 generated labels (See Appendix A.2 for details). On a randomly sampled 100 situations each from H/Bollywood, the model returned 20 samples to be expressing shame in Bollywood and 5 samples in Hollywood. We found poor agreement between human annotators (4 out of 20 samples in Bollywood and 1 out of 5 samples in Hollywood where both annotators agreed with GPT-4). Human annotators across cultures tend to cross-label shame and guilt Troiano et al. (2019) and follow-up discussions with annotators revealed similar tendencies. One such example from Hollywood is provided below:
‘‘about feeding the poor, I’ve never done any of that. God tells us to love everybody. I’ve hated people… my family, my family, ’’
Here, the speaker is probably expressing guilt, shame, or both. Given subjective beliefs and poor agreement with generated outputs, we adopted a high precision-low recall strategy and limited our further analysis to explicit mentions of shame and pride.
For our analysis, we created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood. Cross-cultural differences also exist in speaking styles across cultures and we thus formed a control set of dialogues unrelated to shame and pride for both movie industries to remove the variations in language markers owing to culture-specific speaking styles (See Table 2 for dialogue distribution).
@@ 2.1 Extracting Shame and Pride related Expressions
Table 3: Pearson r for LIWC categories significantly correlated ($p<0.05$) with shame for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values). Note, $1^{st}$ person sing pronouns are strongly correlated with Hollywood-shame whereas the correlation with $2^{nd}$ person pronouns and social references are up to 3 times stronger in Bollywood compared to Hollywood. See Table A7 and A9 for the complete set of correlations.
@@ 3 Approach
Fig. 1 illustrates our study design comprising a blend of two key approaches (a) a psychosocial vocabulary approach to measure how expressions of shame and pride differ and (b) prompting LLM approach to extract reasons behind them and how they differ between India and the U.S.A.
@@ 3.2 Pretrained LLM: GPT-4 Chat
Identifying why shame/pride is expressed in discourse requires multicultural world knowledge. LLMs such as GPT-4 exhibit a superlative pragmatic understanding of the world around us and are increasingly used for extracting implicit meanings and beliefs Pan et al. (2023); Törnberg (2023). We used GPT-4 chat in a two-shot setting ( See Tables A2 and A3 for prompts) to identify:
  • who is $<$being shamed/feeling proud$>$ in the given movie discourse, and what is their gender? and,
@@
Since there are at least two characters in a discourse, the first question orients the LLM to focus on the person experiencing the social sanction or approval and then identify their gender. The output for the second question serves as the implicit social expectation that led to the expression of shame/pride in the culture. Asking for "reason" leading to the expression of social emotion encourages LLMs to retrieve expectations from the provided context, mitigating potential Anglocentric tendencies Havaldar et al. (2023b).
Thematic Analysis of Reasons Behind Shame/Pride-related Expressions
To capture overarching themes in social expectations that led to shame/pride-related expressions in Indian and American societies, we performed agglomerative clustering after embedding unique shame and pride-related reasons using SBERT embeddings Reimers and Gurevych (2019). Unlike movie dialogues, reasons extracted from GPT-4 chat are short phrases (See Table 6) and are devoid of culture-specific language style markers (i.e., Indian English vs American English) thus, no control for culture-specific language is needed.
@@
Bollywood Hollywood
shame pride shame pride
male 2690 1259 591 776
female 1215 326 246 236
Not clear 1416 652 319 719
total 5321 2237 1156 1731
Table 5: Reason and Gender distribution (with duplicates) for dialogues for which GPT-4 predicted male or female. The duplicate reasons are not removed as their frequency reflects their prevalence and is useful for estimating gender association.
@@ 4 Results
4.1 How is shame/pride expressed across cultures?
Shame is associated with negative emotions, power, and morality in both movie industries (See Table 3), however, significant cultural nuances exist in its manifestation. In Hollywood, shame is more I-focused with remorse (e.g., sadness, anxiety) and past-focused language, whereas in Bollywood, it is more You-focused with frequent social references and present-focused language. In Bollywood, psychosocial categories associated with shame reveal a larger social context (note the contrast in family i.e., nuclear family in Hollywood vs extended family in Bollywood) and social situations (i.e., anger, swear, conflict). Categories female and sexual are exclusively correlated with shame in Bollywood potentially indicating an honor system in collectivist societies Caffaro et al. (2014) whereas it is illness that is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
A similar dichotomy is observed for pride (See Table 4). Pride in Hollywood is more centered toward family (note, family in Bollywood is male-centric and comprises extended family (such as uncle)) and self-centered (e.g., I am proud of __) whereas it is achievement/power focused and "We"-centered in Bollywood movies (with othering (us vs they)- See Table A8). Male references are more strongly associated with pride than females in Hollywood whereas the female category has an insignificant ($p>0.05$) correlation with pride in Bollywood. Pride is exclusively positive (see POSEMO in Table A10) in Hollywood whereas in Bollywood, we suppose that it relates to honor more (e.g., protecting family’s pride, bringing pride to family) and is not always used in a positive context.
4.2 How do social expectations behind shame/pride vary across cultures?
Cross-Cultural Shame/Pride Dataset
For the Bollywood set, GPT-4 chat predicted reasons (also implicit social expectations) for 5321 (98.4%) shame-related dialogues out of 5409, and 2237 (74.6%) pride-related dialogues out of 2999. For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805. Upon manual analysis, we found that pride is also used to express affection, specifically toward close family members such as sons - therefore, a lower number of reasons is associated with pride compared to shame. GPT-4 chat predicted a total of 10,445 social expectations (See Table 5 for distribution). Prompting "reasons" behind shame and pride allowed us to capture high specificity in cultural norms (See Tables 6 for shame-related expectations and A6 for pride-related expectations).
4.2.1 Manual Evaluation
Two volunteers manually verified the predicted gender for "the person experiencing shame/pride" and the predicted "reason" in a randomly sampled set of $100$ dialogues each from Hollywood and Bollywood. The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society. Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social expectations of the U.S.A.